Find an equation for the line which passes through (-1,2) and is perpendicular to the line containing (0,2) and (4,6). 


The equation of the line is _______ 
Answer #1

Given points (0,2) and (4,6)

slope(m1) = (6-2)/(4-0) = 1

Perpendicular line slope(m2) = -1/m1 = -1

given point (-1,2)

Equation , y-2 = -1(x+1)

y = - x + 1 . (Answer)